OpenStudy (anonymous):

The formula for the volume (V) of a box is V = lwh. Find h, if V = 0.0144, l = 0.08, and w = 0.6. Answer h = 3 h = 0.03 h = 0.3 h = 0.003

OpenStudy (anonymous):

@ktnguyen1 D?

OpenStudy (anonymous):

h= V/lw = 0.0144/0.08(0.6)

OpenStudy (anonymous):

0.3